HARMONIC ANALYSIS ON THE FINITE

TWISTED POINCARÉ UPPER HALF PLANE

Jorge Vargas Jorge Soto-Andrade

Abstract. We prove that the induced representation from a non trivial character of the Coxeter torus of GL(2, F), for a finite field F, is multiplicity-free; we give an explicit description of the corresponding (twisted) spherical functions and a version of the Heisenberg Uncertainty Principle.

Typeset by AMS-TEX